User Tools

Site Tools



RCP into Chan 1 on Frequency Translator, LCP into Chan 2

Sky Freq 8425, Agilent 13050, Fixed 4100 = IF 525, SMY LO 627

DAS profile = VSOP.PRO

used 10th harmonics for coherence at 841.4 (→8414) and 843.0 (→8430)

Recorded to disks SWIN V013A, labels SWIN_V013A_P1 and SWIN_V013A_P2


Before the experiment we realised that we had a bad connection/cable in the LCP. We tracked down the cable and determined that the problem was a faulty connector (after we replaced it the problem appeared to go away). On the completion of this experiment we did further calibration scans and saw that the problem had persisted. i.e. this experiment will be subject to the side effects of a faulty cable, expect 'dropouts' in the signal.

There also may be an additional problem. During the experiment after this one we found that the oscillator that was used to generate coherence tones was left powered on (tone off) without being locked on the external reference, which may have caused the site-wide reference to become unreliable.This may have effected the sensitivity of Ceduna throughout this experiment. We can not be sure if ext ref was on or off during this experiment but it would be likely that it was off if the sensitivity of Ceduna is lower than expected.

lbaops/lbamar2008/v277acdlog.txt · Last modified: 2015/12/18 16:38 (external edit)